40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 670     Entry time: Tue Jul 15 09:47:09 2008
Author: steve 
Type: Update 
Category: PEM 
Subject: lab temps and particles 
All air condtion units were serviced last Friday.
AC filters are trying to control our particle counts but they have no capacity to match bad Pasadena conditions.
IFO room filters at CES were really clean.
Air make up filters inside and outside were dirty.
They showed the construction effect.
Control room and clean assembly units needed all filters replaced.

Note: the PSL-FSS_RCTEMP droped o.1C when enclouser HEPAs were turned back on
The RC temp controller should be better than that!
Attachment 1: temps24d.jpg  109 kB  | Hide | Hide all
temps24d.jpg
ELOG V3.1.3-